# Hyperband-based Bayesian Optimization for Black-box Prompt Selection

### v1

@@ Abstract
Optimal prompt selection is crucial for maximizing large language model (LLM) performance on downstream tasks. As the most powerful models are proprietary and can only be invoked via an API, users often manually refine prompts in a black-box setting by adjusting instructions and few-shot examples until they achieve good performance as measured on a validation set. Recent methods addressing static black-box prompt selection face significant limitations: They often fail to leverage the inherent structure of prompts, treating instructions and few-shot exemplars as a single block of text. Moreover, they often lack query-efficiency by evaluating prompts on all validation instances, or risk sub-optimal selection of a prompt by using random subsets of validation instances. We introduce HbBoPs, a novel Hyperband-based Bayesian optimization method for black-box prompt selection addressing these key limitations. Our approach combines a structural-aware deep kernel Gaussian Process to model prompt performance with Hyperband as a multi-fidelity scheduler to select the number of validation instances for prompt evaluations. The structural-aware modeling approach utilizes separate embeddings for instructions and few-shot exemplars, enhancing the surrogate model’s ability to capture prompt performance and predict which prompt to evaluate next in a sample-efficient manner. Together with Hyperband as a multi-fidelity scheduler we further enable query-efficiency by adaptively allocating resources across different fidelity levels, keeping the total number of validation instances prompts are evaluated on low. Extensive evaluation across ten benchmarks and three LLMs demonstrate that HbBoPs outperforms state-of-the-art methods.
@@ 1 Introduction
In recent years, pre-trained auto-regressive large language models (LLMs) have demonstrated remarkable capabilities in addressing a wide range of machine learning tasks involving natural language (Brown et al., 2020; Liu et al., 2023), such as Q&A (Joshi et al., 2017; Clark et al., 2018), text summarization (Koupaee & Wang, 2018), text generation (Hendrycks et al., 2020), and mathematical problem solving (Hendrycks et al., 2020; Cobbe et al., 2021). The performance on these tasks relies heavily on prompt engineering, where the input is formatted within a carefully designed prompt that may include instructions, few-shot examples, and additional information, because LLMs are highly sensitive to their input (Zhou et al., 2022; Honovich et al., 2022; Lu et al., 2021; Wu et al., 2024) . As such, selecting the optimal instruction and the examples to include in a few-shot exemplar becomes an imperative task (Liu et al., 2023; Ye et al., 2023).
@@
Static black-box prompt optimization and selection have recently gained much attention (Sun et al., 2022b; Chen et al., 2023; Lin et al., 2023; Wu et al., 2024; Shi et al., 2024). The goal is to construct or identify a single prompt for a black-box LLM that performs well across all instances of a downstream task. This process involves evaluating different prompts on a validation set and using derivative-free techniques to guide optimization or selection. The static black-box setting allows for offline optimization, with the resulting prompt being used for the downstream task.
@@
While much research has focused on automatically generating new prompts (Sun et al., 2022b; Xu et al., 2022; Zhou et al., 2022; Chen et al., 2023; Fernando et al., 2023; Lin et al., 2023), there is growing interest in efficiently selecting prompts from a predefined candidate pool (Shi et al., 2024). This is partly because many optimization techniques involve generating a large candidate pool before identifying the best prompt (Xu et al., 2022; Zhou et al., 2022; Fernando et al., 2023; Prasad et al., 2022). Methods that operate in this selection setting are given by MIPROv2 (Opsahl-Ong et al., 2024), EASE (Wu et al., 2024), TRIPLE-SH and TRIPLE-GSE (Shi et al., 2024).
@@
For these methods we identify the following limitations: (1) With the exception of MIPROv2 they are not explicitly designed for the problem of jointly selecting instructions and few-shot exemplars as they either focus on exemplars (EASE) or instructions (TRIPLE-SH and TRIPLE-GSE). (2) No method is both sample-efficient (allows for evaluating fewer prompts) and query-efficient (reduces the total number of LLM calls by not evaluating all prompts on all validation instances).
@@
In this work, we propose HbBoPs to tackle these limitations. Our main contributions are the following: (1) We design a structural-aware deep kernel Gaussian Process operating on separate embeddings of instructions and few-shot exemplars that learns a lower-dimensional latent representation of prompts in an end-to-end fashion. (2) We propose to use Hyperband as a multi-fidelity scheduler for prompt selection that governs the number of validation instances prompts are evaluated on. (3) We build a novel method, HbBoPs, that relies on this structural-aware deep kernel GP to make a BO proposal within Hyperband allowing for both sample- and query-efficiency. (4) We compare HbBoPs against four baselines and four state-of-the-art algorithms across 10 benchmark tasks and 3 LLMs demonstrating that HbBoPs performs best with respect to identifying a well-performing prompt after a given budget of total LLM calls but also exhibits strongest anytime performance during the selection process. (5) We perform an ablation study of the components of HbBoPs to gain insight into its inner working and further demonstrate its robustness to the choice of the encoder model used to obtain embeddings of prompts.
@@ 2 Problem Statement
Let =$\mathcal{I}=\{i_{1},\ldots,i_{l}\}$ denote a finite set of instructions and =$\mathcal{E}=\{e_{1},\ldots,e_{m}\}$ a finite set of few-shot exemplars. Note that by exemplar we refer to an ordered tuple of a given number of input-output examples of a task. Let $\mathcal{P}=\mathcal{I}\times\mathcal{E}$ be the set of prompts which are generated by combining each $i\in\mathcal{I}$ with each $e\in\mathcal{E}$.
Instructions can be generated either manually by experts or automatically by LLM-based methods, e.g., Automatic Prompt Engineering (APE) by (Zhou et al., 2022). Few-shot exemplars can be generated by selecting different input-output instances from the training set.
A prompt $p\in\mathcal{P}$ is instantiated by combining it with a given input $x\in\mathcal{X}$ for which the LLM, $h:(\mathcal{P}\times\mathcal{X})\rightarrow\mathcal{Y},h([p,x])\mapsto\hat{y}$, produces an output $\hat{y}\in\mathcal{Y}$. We will use $h_{p}$ to denote $h([p,x])$ as a shorthand.
We make no assumptions regarding the nature of the LLM and treat it as a black-box and the LLM returns output given input without any additional information, i.e., no access to model parameters, gradients, or token probabilities.
Based on a validation set $\{(x_{i},y_{i})\}_{i=1}^{n_{\mathrm{valid}}}$, evaluating a prompt is performed by comparing the ground truth output $y_{i}$ to the output $\hat{y}_{i}=h_{p}(x_{i})$ generated by the LLM based on a point-wise loss function $l:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R},(y,\hat{y})\mapsto l(y,\hat{y})$. This loss function quantifies how close the output generated by the LLM is to the ground truth. For example, a loss function based on the widely used exact match (Chang et al., 2024) scoring function is given by:
@@
Our task is to identify a single prompt $p\in\mathcal{P}$ which performs optimal with respect to the loss in expectation:
@@
Here the expectation is taken over all input-output instances $(x,y)$. In practice Equation (2) is approximated based on the validation instances available:
$f(p)\coloneqq\frac{1}{n_{\mathrm{valid}}}\sum_{i=1}^{n_{\mathrm{valid}}}l(y_{i},h_{p}(x_{i}))$ (3)
We refer to this setting as the static setting as we are searching for a single optimal prompt on the target downstream task. Note that due to the non-deterministic nature of LLMs, $f$ itself can in general only be observed with noise. We will denote by
@@
the observed validation error of the LLM configured to use prompt $p$. $f$ is a black-box function as the LLM is a black-box and no analytic description or gradient information is available.
Given the finite search space $\mathcal{P}$, our goal is to identify the best prompt using as few LLM calls as possible.
@@ 3 Method
We want to learn a surrogate model of the black-box function $f$ described in Equation (3) that models the LLM validation error of each prompt on the downstream task. This surrogate is used to predict the validation error of the not-evaluated prompts during the optimization process (the selection of the next candidate prompt). As discussed in Section 2, prompts are not merely blocks of text but possess a specific structure composed of various building blocks. While it may seem intuitive to embed prompts and train a Gaussian Process (GP) on high-dimensional embeddings, we argue that this simplistic approach fails to adequately capture the intricate structure of prompts. To address this limitation, we propose a structure-aware deep kernel (Section 3.2).
3.1 Gaussian Process as a Surrogate Model
@@
To learn a surrogate model, we collect, for a given optimization step $t$, design data $\mathcal{D}_{t}:=\{(p_{j},v_{j})\}_{j=1}^{t}$, where each tuple is given by the prompt $p_{j}\in\mathcal{P}$ and its validation error $v_{j}$ described in Equation (4) at the $j$-th previous optimization step. This design data is the collection of prompts and their performance during the sequential optimization process. To learn a model that maps prompts to their validation errors, we embed each prompt into a $d$-dimensional numeric space making use of pre-trained language encoders. Let $enc:\mathcal{P}\rightarrow\mathbb{R}^{d},p\mapsto\mathbf{z}$ be the encoding function, then we augment the design data ($\mathcal{D}_{t}=\{(p_{j},\mathbf{z}_{j},v_{j})\}_{j=1}^{t}$) with $\mathbf{z}_{j}$, the embedding of the prompt $p_{j}$.
@@
We want to use a GP as our surrogate model since it allows for flexible probabilistic modeling of black-box functions by returning a point estimate and well-calibrated uncertainty estimations in the form of a Gaussian posterior predictive distribution (Williams & Rasmussen, 2006). In the following, we assume a GP prior over $f$ in the $d$-dimensional space of the embedded prompts, $f(\mathbf{z})\sim\mathcal{GP}(m,k)$; $\mathbf{f}\sim\mathcal{N}(m(\mathbf{Z}),k(\mathbf{Z},\mathbf{Z}|\mathbf{\theta}))$, where $m$ is the prior mean function usually set to $\mathbf{0}$ (zero mean), $k$ is the covariance function depending on kernel parameters $\mathbf{\theta}$, and $\mathbf{Z}$ is a matrix of prompt embeddings.
@@
Then, given the design data $\mathcal{D}_{t}$, and new prompts $\mathbf{p}_{\star}$ with their embeddings $\mathbf{Z}_{\star}$, the function $\mathbf{f}_{\star}$ is modelled as a random variable that is jointly Gaussian distributed with all previously observed validation errors $\mathbf{v}=(v_{1},\ldots,v_{t})$. In short
@@
$\begin{bmatrix}\mathbf{v}\\ \mathbf{f}_{\star}\end{bmatrix}\sim\mathcal{N}\left(m(\mathbf{Z},\mathbf{Z}_{\star}),\begin{bmatrix}\mathbf{K}_{t}&\mathbf{K}_{\star}\\ \mathbf{K}^{\intercal}_{\star}&\mathbf{K}_{\star\star},\end{bmatrix}\right)$
@@
$\begin{split}\mathbb{E}[\mathbf{f}_{\star}|\mathbf{Z},\mathbf{v},\mathbf{Z}_{\star}]=\mathbf{K}^{\intercal}_{\star}(\mathbf{K}_{t})^{-1}\mathbf{v},\\ \mathrm{cov}[\mathbf{f}_{\star}|\mathbf{Z},\mathbf{Z}_{\star}]=\mathbf{K}_{\star\star}-\mathbf{K}^{\intercal}_{\star}(\mathbf{K}_{t})^{-1}\mathbf{K}_{\star},\end{split}$ (5)
where common choices for kernel functions $k$ are given by the squared exponential kernel or variations of the Matérn kernel (see, e.g., Williams & Rasmussen, 2006, Chapter 4).
At this point, we could proceed and train a ”vanilla” GP as outlined above on the $d$-dimensional space of embedded prompts. However, as stated in many previous works on BO (Kandasamy et al., 2015; Wang et al., 2016; Gardner et al., 2017; Eriksson et al., 2019; Eriksson & Jankowiak, 2021), GPs struggle with high-dimensional input as the one that we have in our design data $\mathcal{D}_{t}$, e.g., the dimension of the BERT [CLS] token embedding is 768. Moreover, dimensionality reduction techniques such as principle component analysis (PCA) or random projections could be used, however, such techniques are unsupervised and will not result in a lower-dimensional representation that is aligned with the downstream performance of prompts (see also Figure 3 in Appendix A for an illustration). Additionally, using a single embedding of the whole prompt treated as a block of text ignores that the prompt is composed of different building blocks carrying distinct information. Below, we present our solution to this problem.

### v2

@@ Abstract
Optimal prompt selection is crucial for maximizing large language model (LLM) performance on downstream tasks, especially in black-box settings where models are only accessible via APIs. Black-box prompt selection is challenging due to potentially large, combinatorial search spaces, absence of gradient information, and high evaluation cost of prompts on a validation set. We propose HbBoPs, a novel method that combines a structural-aware deep kernel Gaussian Process with Hyperband as a multi-fidelity scheduler to efficiently select prompts. HbBoPs uses embeddings of instructions and few-shot exemplars, treating them as modular components within prompts. This enhances the surrogate model’s ability to predict which prompt to evaluate next in a sample-efficient manner. Hyperband improves query-efficiency by adaptively allocating resources across different fidelity levels, reducing the number of validation instances required for evaluating prompts. Extensive experiments across ten diverse benchmarks and three LLMs demonstrate that HbBoPs outperforms state-of-the-art methods in both performance and efficiency.
@@ 1 Introduction
In recent years, pre-trained auto-regressive large language models (LLMs) have demonstrated remarkable capabilities in addressing a wide range of machine learning tasks involving natural language (Brown et al., 2020; Liu et al., 2023), such as Q&A (Joshi et al., 2017; Clark et al., 2018), text summarization (Koupaee & Wang, 2018), text generation (Hendrycks et al., 2021), and mathematical problem-solving (Hendrycks et al., 2021; Cobbe et al., 2021). As LLMs are highly sensitive to their input (Zhou et al., 2023; Honovich et al., 2023; Lu et al., 2022; Liu et al., 2023; Ye et al., 2023; Wu et al., 2024), performance on these tasks relies on prompt engineering, where the input is formatted within a carefully designed prompt that may include an instruction, a few-shot exemplar, and additional information.
@@
The goal of static black-box prompt optimization and selection (Sun et al., 2022b; Chen et al., 2024; Lin et al., 2024; Wu et al., 2024; Shi et al., 2024) is to construct or identify a single prompt for a black-box LLM that, in expectation, performs well across all instances of a downstream task. This process involves evaluating different prompts on a validation set and using derivative-free techniques to guide optimization or selection. The static black-box setting allows for offline optimization, with the resulting optimal prompt being used for the downstream task.
@@
While much research has focused on automatically generating new prompts (Sun et al., 2022b; Xu et al., 2022; Zhou et al., 2023; Chen et al., 2024; Fernando et al., 2024; Lin et al., 2024), there is growing interest in efficiently selecting prompts from a predefined pool of candidates (Shi et al., 2024). This is because many prompt optimization techniques involve generating a large candidate pool a priori before identifying the best prompt (Xu et al., 2022; Zhou et al., 2023; Fernando et al., 2024; Prasad et al., 2023). Moreover, recent empirical findings indicate that few-shot exemplars often contribute more strongly to prompt performance than the instruction of a prompt itself, with the best results typically achieved through the optimal selection of both components (Wan et al., 2024).
The task of black-box prompt selection is challenging. First, the search space can be extensive, as instructions and few-shot exemplars form a combinatorial set of candidates. Second, black-box LLMs make it impossible to directly optimize based on gradient information. Third, evaluating a prompt is time-consuming and costly, as each evaluation involves querying the LLM on multiple validation instances of a task. This calls for the development of selection methods that can efficiently explore the space of candidate prompts while keeping the total number of LLM calls low.
@@
Existing methods for prompt selection in this setting include MIPROv2 (Opsahl-Ong et al., 2024), EASE (Wu et al., 2024), TRIPLE-SH, and TRIPLE-GSE (Shi et al., 2024). We identify the following limitations among these approaches: (1) Except for MIPROv2, these methods are not explicitly designed to address the problem of jointly selecting instructions and few-shot exemplars. Specifically, EASE primarily focuses on exemplar selection, while TRIPLE-SH and TRIPLE-GSE focus on instruction selection. Although EASE can be applied to joint selection by treating the entire prompt as an unstructured block of text, it does not exploit the compositional structure of prompts. (2) No method is both sample-efficient (allowing for evaluating fewer prompts by relying on a surrogate model) and query-efficient (reducing the total number of LLM calls by not evaluating prompts on all available validation instances).
@@
In this work, we propose HbBoPs (Hyperband-based Bayesian optimization for black-box Prompt selection) addressing these limitations. Our main contributions are the following: (1) We introduce a structural-aware deep kernel Gaussian Process (GP) that learns a low-dimensional prompt representation from separate embeddings of instructions and few-shot exemplars in an end-to-end fashion to predict prompt performance. (2) We adopt Hyperband (Li et al., 2018) as a multi-fidelity scheduler for prompt selection that governs the number of validation instances prompts are evaluated on. (3) We introduce a novel method, HbBoPs, that relies on our structural-aware deep kernel GP to make a Bayesian Optimization (BO) proposal within Hyperband and as a result is both sample- and query-efficient. (4) We compare HbBoPs against four baselines and four state-of-the-art methods across ten benchmarks and three LLMs, demonstrating that HbBoPs performs best in identifying a well-performing prompt after a given budget of total LLM calls but also exhibits strongest anytime performance during the selection process. (5) We perform an ablation study of the components of HbBoPs gaining insight into their inner workings and further demonstrate its robustness to the choice of the encoder model used to obtain embeddings.
@@ 2 Problem Statement
Let =$\mathcal{I}=\{i_{1},\ldots,i_{l}\}$ denote a finite set of instructions (task descriptions) and =$\mathcal{E}=\{e_{1},\ldots,e_{m}\}$ a finite set of few-shot exemplars. Note that by exemplar we refer to an ordered tuple of a given number of input-output examples of a task. Let $\mathcal{P}=\mathcal{I}\times\mathcal{E}$ be the set of prompts that are generated by combining each $i\in\mathcal{I}$ with each $e\in\mathcal{E}$.
Instructions can be generated either manually by experts or automatically by LLM-based methods, e.g., Automatic Prompt Engineering (APE; Zhou et al. 2023). Few-shot exemplars can be generated by selecting different input-output instances from a training set of the task.
A prompt $p\in\mathcal{P}$ is instantiated by combining it with a given task input or query $x\in\mathcal{X}$ for which the LLM, $h:(\mathcal{P}\times\mathcal{X})\rightarrow\mathcal{Y},h([p,x])\mapsto\hat{y}$, produces an output $\hat{y}\in\mathcal{Y}$. We will use $h_{p}(x)$ to denote $h([p,x])$ as a shorthand.
We make no assumptions regarding the nature of the LLM and treat it as a black-box. The LLM returns output given input without any additional information, i.e., no access to model parameters, gradients, or token probabilities.
Having access to a validation set $\{(x_{i},y_{i})\}_{i=1}^{n_{\mathrm{valid}}}$, evaluating a prompt is performed by comparing the ground truth output $y_{i}$ to the output $\hat{y}_{i}=h_{p}(x_{i})$ generated by the LLM based on a pointwise loss function $l:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R},(y,\hat{y})\mapsto l(y,\hat{y})$. This loss function quantifies how close the output generated by the LLM is to the ground truth. For example, a loss function based on the widely used exact match (Chang et al., 2024) scoring function is given by:
@@
Our task is to identify a single prompt $p\in\mathcal{P}$ that is optimal with respect to the loss in expectation:
@@
Here, the expectation is taken over all input-output instances $(x,y)$. In practice, however, Equation (2) can only be approximated based on the validation instances available:
$f(p)\coloneqq\frac{1}{n_{\mathrm{valid}}}\sum_{i=1}^{n_{\mathrm{valid}}}l(y_{i},h_{p}(x_{i})).$ (3)
We refer to this setting as the static setting, as we are searching offline for a single optimal prompt on the target downstream task. Note that due to the non-deterministic nature of LLMs (depending on temperature), $f$ itself can in general only be observed with noise.
We will denote by
@@
the observed validation error of the LLM configured to use prompt $p$  [^1] We model the observation noise as homoscedastic Gaussian for analytical convenience.. $f$ is a black-box as the LLM is a black-box and no analytic description or gradient information is available.
The core challenge of static black-box prompt selection is balancing exploration (searching the prompt space $\mathcal{P}$) and efficiency (minimizing costly LLM evaluations). Our goal is to identify the best prompt using as few LLM calls for evaluation as possible.
@@ 3 Method
We want to learn a surrogate model of the black-box function $f$ in Equation (3) that predicts the validation error of prompts on a downstream task based on observed data collected during optimization. This surrogate is used to predict the validation error of the unevaluated prompts during the optimization process. We first describe how GPs can serve as surrogate models in black-box prompt selection (Section 3.1), but highlight the limitations of vanilla GPs on raw, high-dimensional prompt embeddings. To address this, we introduce a structural-aware deep kernel GP (Section 3.2) that relies on structural information of prompts via separate embeddings of their building blocks. We then adopt Hyperband (Li et al., 2018) for query-efficient multi-fidelity prompt selection (Section 3.3) and integrate our structural-aware deep kernel GP via a BO proposal, which results in our final HbBoPs algorithm (Section 3.4).
3.1 Gaussian Process as Surrogate Model
@@
To learn a surrogate model, we collect, at each optimization step $t$, a set of design data $\mathcal{D}_{t}:=\{(p_{j},v_{j})\}_{j=1}^{t}$, where each tuple consists of a prompt $p_{j}\in\mathcal{P}$ and its corresponding validation error $v_{j}$, as defined in Equation (4), recorded at the $j$-th previous step. This design data captures the history of prompts evaluated and their observed performance throughout the sequential optimization process. To learn a model that maps prompts to their validation errors, we embed each prompt into a $d$-dimensional numeric space, making use of pre-trained language encoders. Let $enc:\mathcal{P}\rightarrow\mathbb{R}^{d},p\mapsto\mathbf{z}$ be the encoding function. We then augment the design data, $\mathcal{D}_{t}=\{(p_{j},\mathbf{z}_{j},v_{j})\}_{j=1}^{t}$, where $\mathbf{z}_{j}$ is the embedding of prompt $p_{j}$. Since we are concerned with black-box prompt selection, it is natural to use embeddings as feature representations of prompts. Notably, recent work (Tang et al., 2024) has shown that (LLM) embeddings can in general serve as effective features for high-dimensional regression tasks, even in domains where the input data is not textual.
@@
We want to use a GP as a surrogate model since it allows for flexible probabilistic modeling of black-box functions by returning a point estimate and well-calibrated uncertainty estimates in the form of a Gaussian posterior predictive distribution (Williams & Rasmussen, 2006). In the following, we assume a GP prior over $f$ in the $d$-dimensional space of embedded prompts, $f(\mathbf{z})\sim\mathcal{GP}(m,k)$; $\mathbf{f}\sim\mathcal{N}(m(\mathbf{Z}),k(\mathbf{Z},\mathbf{Z}|\mathbf{\theta}))$, where $m$ is the prior mean function, usually set to zero, $k$ is the covariance function depending on kernel parameters $\mathbf{\theta}$, and $\mathbf{Z}$ is a matrix of prompt embeddings.
@@
Given the design data $\mathcal{D}_{t}$ and new prompts $\mathbf{p}_{\star}$ with their embeddings $\mathbf{Z}_{\star}$, the function $\mathbf{f}_{\star}$ is modeled as a random variable that is jointly Gaussian distributed with all previously observed validation errors $\mathbf{v}=(v_{1},\ldots,v_{t})$.
In short, this can be written as
@@
$\begin{bmatrix}\mathbf{v}\\ \mathbf{f}_{\star}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}m(\mathbf{Z})\ m(\mathbf{Z}_{\star})\end{bmatrix},\begin{bmatrix}\mathbf{K}_{t}&\mathbf{K}_{\star}\\ \mathbf{K}^{\intercal}_{\star}&\mathbf{K}_{\star\star},\end{bmatrix}\right),$
@@
$\begin{split}\mathbb{E}[\mathbf{f}_{\star}|\mathbf{Z},\mathbf{v},\mathbf{Z}_{\star}]&=\mathbf{K}^{\intercal}_{\star}(\mathbf{K}_{t})^{-1}\mathbf{v},\\ \mathrm{cov}[\mathbf{f}_{\star}|\mathbf{Z},\mathbf{Z}_{\star}]&=\mathbf{K}_{\star\star}-\mathbf{K}^{\intercal}_{\star}(\mathbf{K}_{t})^{-1}\mathbf{K}_{\star},\end{split}$ (5)
where common choices for the kernel function $k$ are given by the squared exponential kernel or variations of the Matérn kernel (see, e.g., Williams & Rasmussen, 2006, Chapter 4).
At this point, we could proceed and train a vanilla GP as outlined above on the $d$-dimensional space of embedded prompts. However, as stated in many previous works on BO (Kandasamy et al., 2015; Wang et al., 2016; Gardner et al., 2017; Eriksson et al., 2019; Eriksson & Jankowiak, 2021), GPs struggle with high-dimensional input such as that found in our design data $\mathcal{D}_{t}$, e.g., the dimensionality of BERT’s (Devlin et al., 2019) [CLS] token embedding is already 768. In general, dimensionality reduction techniques such as principal component analysis (PCA) or random projections could be used. However, such techniques are unsupervised and will not yield a lower-dimensional representation that is aligned with the downstream performance of prompts (see also Figure 3 in Appendix A for an illustration). Additionally, using a single embedding of the whole prompt treated as a block of text ignores that the prompt is composed of different building blocks with distinct structural information. Below, we present our solution.